OpenStudy (anonymous):

The tiles E, T, F, U, N, X, and P of a word game are placed face down in the lid of the game. If two tiles are chosen at random, find each probability. 1. P(E, then N), if replacement occurs 2. P(2 consonants), if replacement occurs 3. P(T, then D), if replacement occurs 4. P(2 consonants), if no replacement occurs 5. P(selecting the same letter twice), if no replacement occurs

OpenStudy (anonymous):

1) and 3) the same answer (1 )/(7 ) * (1 )/(7 ) = 1/(49 )? 4) I don't need any more

OpenStudy (anonymous):

1st one \[\frac{1}{7}\times \frac{1}{7}\]

OpenStudy (anonymous):

same as 3

OpenStudy (anonymous):

second one \[(\frac{5}{7})^2\]

OpenStudy (anonymous):

5) probability is zero

OpenStudy (anonymous):

since it is not possible to pick the same letter twice if you don't replace them